OpenStudy (anonymous):

A company sells boxes of small blue pom-poms (b) for $42 and boxes of large orange pom-poms (p) for $52. They make batches of small blue pom-poms that fill 8 boxes and batches of large orange pom-poms that fill 4 boxes. The company only has 40 boxes. Which system of equations helps the company plan to make $350?

OpenStudy (anonymous):

b = 40 - p 8b + 4p = 40 b = 40 - p 42b + 52p = 350 8b = 40 - 4p 42b - 52p = 350 8b = 40 - 4p 42b + 52p = 350